Cash Payback Period, Net Present Value Analysis, and Qualitative Considerations

The plant manager of Shenzhen Electronics Company is considering the purchase of new automated assembly equipment. The new equipment will cost $175,000. The manager believes that the new investment will result in direct labor savings of $35,000 per year for 10 years.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. What is the payback period on this project?
years

b. What is the net present value, assuming a 12% rate of return? Use the table provided above. Round to the nearest whole dollar.

Net present value $
Answer #1

Answers:

Payback Period = 5 Years

Net present Value = 22,750

Payback Period = Investment / Annual Cash Flows = 175,000 /35,000 =5 years

Net Present Value = Present value of Annual Cash flows - Investment

Discount Rate = 12%

= 35,000*5.650-175,000

= 197,750-175,000 =22,750 (answer)